OpenStudy (anonymous):

Find the length and width of a rectangle, given that width is one-half of the length and the perimeter of the rectangle is 54 cm.

OpenStudy (cggurumanjunath):

w=0.5l p=2(l+w)=54

OpenStudy (anonymous):

so whats the answer

OpenStudy (cggurumanjunath):

u find it @lubavasilyuk

OpenStudy (anonymous):

i dont get how

OpenStudy (cggurumanjunath):

w=0.5l put w=0.5l in p=2(l+w)

OpenStudy (anonymous):

p=2(l+0.51)

OpenStudy (cggurumanjunath):

good

OpenStudy (anonymous):

wait the width is 51 correct?

OpenStudy (cggurumanjunath):

p=2(l+0.51)=54

OpenStudy (cggurumanjunath):

p=2(l+w)

OpenStudy (cggurumanjunath):

here p=perimeter

OpenStudy (cggurumanjunath):

l=length

OpenStudy (cggurumanjunath):

w=width

OpenStudy (cggurumanjunath):

p=2(l+0.51)=54

OpenStudy (anonymous):

length is 3?

OpenStudy (cggurumanjunath):

so 2(l+0.5l)=54

OpenStudy (cggurumanjunath):

2(1.5l)=54

OpenStudy (cggurumanjunath):

3l=54

OpenStudy (cggurumanjunath):

l=54/3

OpenStudy (cggurumanjunath):

@lubavasilyuk

OpenStudy (anonymous):

18

OpenStudy (cggurumanjunath):

:)

OpenStudy (anonymous):

so the width is 18?

OpenStudy (cggurumanjunath):

w=0.5l

OpenStudy (cggurumanjunath):

bingo

OpenStudy (anonymous):

oh lenght i meant

OpenStudy (cggurumanjunath):

u got it .

OpenStudy (cggurumanjunath):

:)

OpenStudy (anonymous):

so just to clarify can u type up all that we discuseed like the equation and leght and width?

OpenStudy (cggurumanjunath):

length =18 width=0.5*l=0.5*18=9

OpenStudy (cggurumanjunath):

length =18 width=9

OpenStudy (cggurumanjunath):

is it clear @lubavasilyuk ?

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

and r u 100 percent sure about the answers?

OpenStudy (cggurumanjunath):

width=0.5*length =1/2 *length perimeter=2(length+width)=54cm

OpenStudy (cggurumanjunath):

are u 100 % sure u understood the solution ?